# Denoising diffusion probabilistic models are optimally adaptive to unknown low dimensionality

### v1

@@ 4.2 Convergence guarantees for DDPM
Our convergence theory is stated below, whose proof is deferred to Section 5.2.
@@
Assuming the data lies on a low-dimensional manifold, this concurrent work establishes an iteration complexity that scales linearly in the manifold dimension. While our results and theirs clearly share similarity, we point out several differences here. Firstly, in Potaptchik et al., (2024), the data is assumed to be strictly supported on a constant-scale manifold, and the density function (w.r.t. to the standard volume measure) needs to be both upper and lower bounded uniformly. In comparison, we consider a more general setting where the data support is only required to stay close to a low-dimensional structure whose size can grow polynomially in $k$; also, no boundedness requirement is imposed on the density function. In addition, our paper makes apparent an intimate connection between the DDPM sampler and a properly discretized SDE, which helps explain the rationale behind the adaptivity phenomenon. Finally, our proofs invoke the covering argument to accommodate broader low-dimensional structure, which differs from the techniques in Potaptchik et al., (2024).
5 Main analysis
5.1 Proof of Theorem 1
Step 1: Solving SDE (11b).
Step 2: Establishing the equivalence between (23) and DDPM.
By defining
$\gamma_{n}=e^{-(T-t_{n})},$
we can explicitly calculate the integral as follows
Substitution into equation (24) then gives
5.2 Proof of Theorem 2
We now present the proof of main convergence theory in Theorem 2. To control the KL divergence between the DDPM-type dynamics and our target distribution, a common approach is to connect the KL error with the $\ell_{2}$ discretization error by means of Girsanov’s theorem (e.g., Chen et al., 2022b ; Chen et al., 2023b ; Wu et al., (2024); Benton et al., (2024)). We shall follow this approach in this proof, while in the meantime taking into account the intrinsic low-dimensional structure of the target distribution. In particular, we shall invoke the covering argument to treat the low-dimensional structure approximately as the union of linear subspaces. In particular, As it turns out, the evolution of the posterior mean function, i.e., $\mu_{t}(\cdot)$, can be well calibrated along the forward process with high probability under this framework, which plays a key role in enabling our non-asymptotic convergence analysis.
Our main proof is composed of three steps, supported by a sequence of auxiliary lemmas whose proofs are deferred to Appendix A-B. Throughout the proof, we adopt the following convenient notation:
@@ 5.2 Proof of Theorem 2
In this step, we list several lemmas that help connect the KL error of interest with the mean squared difference between the true and estimated posterior means; these results are now standard for analyzing DDPM-type samplers.
@@
To cope with the term $T_{1}$, we find it helpful to first establish the following result on the coefficients $\{\eta_{t}\}$, whose proof is provided in Appendix A.1.
@@
We now turn to the term $T_{2}$ in (26). Here, $(Y_{t})_{t\in[0,T]}$ stands for the solution to the backward SDE (5) when initialized at $Y_{0}\sim X_{T}$, which satisfies $X_{t}\overset{\mathrm{d}}{=}Y_{T-t}$ (cf. (6)). Towards this end, we intend to first analyze each term in the integral. Define
@@
Towards this end, we begin with the following lemma, which leverages Itô’s isometry to derive an alternative expression for $D_{s,t}$ that lends it well for subsequent analysis. The proof is provided in Appendix A.2.
@@
We are now ready to present a key lemma that can assist in upper bounding $D_{s,t}$; the proof can be found in Appendix A.3.
@@ 6 Discussion
This paper has made progress towards demystifying the unreasonable effectiveness of then DDPM when the target distribution has a low-dimensional structure. We have sharpened the past result by Li and Yan, 2024a , by proving that the iteration complexity scales linearly with the intrinsic dimension $k$. In some sense, our results reveal that the DDPM sampler is optimally adaptive to unknown low dimensionlity, provided that a suitable discretization schedule is adopted. Our result is derived by exploiting the connection between the DDPM sampler and an adaptively discretized SDE, where the non-linear component of the drift time of this SDE is inherently low-dimensional.
We conclude this paper by pointing out several directions worthy of future investigation. To begin with, it remains unclear how the DDPM sampler adapts to unknown low dimensionality when other metrics (e.g., the total-variation distance or the Wasserstein distance) are used to measure distributional discrepancy. Another interesting direction is to examine the adaptivity of other mainstream samplers (e.g., the probability flow ODE (Song et al.,, 2021; Chen et al., 2023d, ; Li et al., 2024c, ; Gao and Zhu,, 2024; Huang et al.,, 2024)) in the face of unknown low-dimensional structure. Moreover, this paper treats the training procedure as a black box and directly assumes a mean squared score estimation error bound. For distributions with low intrinsic dimensions, a desirable score learning algorithm woulclud adapt to the low-dimensional structure too, which has been the focus of a recent line of work (Kadkhodaie et al.,, 2023; Tang and Yang,, 2024; Chen et al., 2024b, ; Azangulov et al.,, 2024; Chen et al., 2023b, ; Wang et al.,, 2024). It would thus be interesting to develop an end-to-end theory to accommodate both the pre-training phase and the sampling process. Finally, our analysis, which highlights the crucial role of proper parameterization and discretization, might shed light on how to design the coefficients for other algorithms in order to accommodate inherently structured data distributions.
@@ Acknowledgements
Y. Wei is supported in part by the NSF grant CCF-2418156 and the CAREER award DMS-2143215. Y. Chen is supported in part by the Alfred P. Sloan Research Fellowship, the AFOSR grant FA9550-22-1-0198, the ONR grant N00014-22-1-2354, and the NSF grant CCF-2221009. The authors would like to thank Jiadong Liang and Yuchen Wu for their inspiring discussions.
Appendix A Proofs of the main lemmas
A.1 Proof of Lemma 6
@@ A.3 Proof of Lemma 8
where $C_{\mathsf{cover}}^{\prime}$ is a universal constant chosen as $C_{\mathsf{cover}}^{\prime}=C_{0}\cdot C_{\mathsf{cover}}$. Let $\{x_{i}^{*}\}_{1\leq i\leq s}$ be an $\varepsilon_{0}$-net of the data support $\mathcal{X}_{\mathsf{data}}$ (see the definition of an epsilon-net in Wainwright, (2019)), and let $\{\mathcal{B}_{i}\}_{1\leq i\leq s}$ be the corresponding $\varepsilon_{0}$-covering of $\mathcal{X}_{\mathsf{data}}$ obeying $x_{i}^{*}\in\mathcal{B}_{i}$ for all $1\leq i\leq s$. Without loss of generality, we choose the $\varepsilon$-covering in a way such that $\mathcal{B}_{i}$ and $\mathcal{B}_{j}$ are disjoint for all $1\leq i\neq j\leq s$, and $B(x_{i}^{*},\varepsilon_{0}/2)\subset\mathcal{B}_{i}$ for all $1\leq i\leq s$.
@@
$\mathcal{G}$$\coloneqq\big{\{}z\in\mathbb{R}^{d}\mid\|z\|_{2}\leq 2\sqrt{d}+\sqrt{C_{6}k\log k},\quad\text{and}$
@@
where $C_{6}>0$ is some sufficiently large constant. The index set $\mathcal{I}$ can be regarded as the collection of sets in $\varepsilon_{0}$-covering with large enough probability, and we can see later that $\mathcal{G}$ is a high-probability region of a standard Gaussian random vector. Further, define a high-probability region of the random vector $X_{t}$ as:
@@
whose high-probability property is confirmed by the following lemma. The proof of this lemma is provided in Appendix B.2.
@@
To obtain tight control of $\mathbb{E}\left[\mathrm{Tr}\big{(}\Sigma_{t}(X_{t})\big{)}\right]$, we look at the events $\{X_{t}\in\mathcal{T}_{t}\}$ and $\{X_{t}\notin\mathcal{T}_{t}\}$ separately.
@@
The following lemma characterizes the typical performance of $X_{0}$, whose proof can be found in Appendix B.3.
@@
Consider any $x\in\mathcal{T}_{t}$, and suppose $x=\sqrt{1-\sigma_{t}^{2}}x_{0}^{\prime}+\sigma_{t}z$ with $x_{0}^{\prime}\in\mathcal{B}_{i(x)}$ and $z\in\mathcal{G}$. We have
@@
holds for all $C\geq C_{7}$ provided $C_{7}>0$ is some large enough constant.
Armed with the sub-Gaussian tail probability bound in Lemma 12, one can directly check that, for all $x\in\mathcal{T}_{t}$,
$\mathrm{Tr}\big{(}\Sigma_{t}(x)\big{)}\leq\mathbb{E}\big{[}\big{\|}X_{0}-x_{i(x)}^{*}\big{\|}_{2}^{2}\mid X_{t}=x\big{]}\lesssim\frac{\sigma_{t}^{2}}{1-\sigma_{t}^{2}}k\log k,$ (47)
with the proviso that $\sigma_{t}^{2}/(1-\sigma_{t}^{2})\geq\sigma_{\delta}^{2}=\Omega(1/\mathrm{poly}(k))$.
@@
$\mathrm{Tr}\left(\Sigma_{t}(x)\right)\leq\mathbb{E}[\|X_{0}\|_{2}^{2}\mid X_{t}=x]\leq\sup_{x\in\mathcal{X}_{\mathsf{data}}}\|x\|_{2}^{2}\leq k^{2C_{R}}.$ (48)
@@
provided that $C_{6}$ is sufficiently large and $\sigma_{t}^{2}/(1-\sigma_{t}^{2})=\Omega(1/\mathrm{poly}(k))$. Inequality (49) taken together with inequality (44) concludes the proof.
Appendix B Proofs of the auxiliary lemmas
B.1 Proof of Lemma 1
@@
$N^{\mathsf{cover}}(\mathcal{M},\|\cdot\|_{2},\varepsilon_{0})\leq\frac{\mathrm{vol}(\mathcal{M})}{\omega_{k}}\left(\frac{\pi}{2}\right)^{k}\varepsilon_{0}^{-k},$
@@
$N^{\mathsf{cover}}(\mathcal{M}^{\varepsilon_{0}},\|\cdot\|_{2},\varepsilon_{0})\leq N^{\mathsf{cover}}(\mathcal{M},\|\cdot\|_{2},\varepsilon_{0}/2)\leq\frac{\mathrm{vol}(\mathcal{M})}{\omega_{k}}\pi^{k}\varepsilon_{0}^{-k}.$
@@
$\log N^{\mathsf{cover}}(\mathcal{M}^{\varepsilon_{0}},\|\cdot\|_{2},\varepsilon_{0})$ $\leq k\log\varepsilon_{0}^{-1}+\log\mathrm{vol}(\mathcal{M})-\log\omega_{k}+k\log\pi$
$\leq k\left(\log\varepsilon_{0}^{-1}+(\mathrm{vol}(\mathcal{M}))^{1/k}+C_{4}\right),$
where $C_{4}$ is some numerical constant. Therefore, if $\mathcal{M}$ satisfies $\tau_{\mathcal{M}}=\Omega(k^{-C_{1}})$ and $\mathrm{vol}(\mathcal{M})=O(k^{C_{2}})$ for some large enough constants $C_{1},C_{2}>0$, we arrive at
@@
B.2 Proof of Lemma 11
Since $\mathcal{T}_{t}$ is characterized by two components $\bigcup_{i\in\mathcal{I}}\mathcal{B}_{i}$ and $\mathcal{G}$, we analyze them separately. Let us begin with the set $\bigcup_{i\in\mathcal{I}}\mathcal{B}_{i}$, which satisfies
$\mathbb{P}\bigg{(}X_{0}\notin\bigcup_{i\in\mathcal{I}}\mathcal{B}_{i}\bigg{)}$ $\overset{(a)}{\leq}\sum_{i\in\mathcal{I}}\exp(-C_{6}k\log k)$

### v2

@@ 4.2 Convergence guarantees for DDPM
Our convergence theory is stated below, whose proof is deferred to Section 5.
@@
Assuming the data lies on a low-dimensional manifold, this concurrent work, posted two weeks before ours, established an iteration complexity that scales linearly in the manifold dimension. As alluded to previously, our results and theirs share clearly similarities in both the aims and the results. Here, let us also point out several differences here. In Potaptchik et al., (2024), the data is assumed to be strictly supported on a constant-scale manifold with diameter less than $1$ (see Assumptions A and B therein), and the density function (w.r.t. to the standard volume measure) needs to be both upper and lower bounded uniformly (see Assumption C therein). In comparison, we consider a more general setting where the data support is only required to stay close to a low-dimensional structure whose size can grow polynomially in $k$. Also, no boundedness requirement is imposed on the density function in our theory, thereby accommodating a much broader scenario. Notably, the theory in Potaptchik et al., (2024) only applies to data with $\mathbb{E}\big{[}\|X_{0}\|_{2}^{2}\big{]}=O(1)$ second moment (as a consequence of their assumption that the diameter of the manifold is $O(1)$), while ours is able to accommodating polynomia lly large second moment $\mathbb{E}\big{[}\|X_{0}\|_{2}^{2}\big{]}=\mathrm{poly}(k)$. Our proof invokes the covering-style argument to accommodate broader low-dimensional structure, which differs from the techniques in Potaptchik et al., (2024).
5 Main analysis (Proof of Theorem 2)
We now present the proof of main convergence theory in Theorem 2. To control the KL divergence between the DDPM-type dynamics and our target distribution, a common approach is to connect the KL error with the $\ell_{2}$ discretization error by means of Girsanov’s theorem (e.g., Chen et al., 2022b ; Chen et al., 2023b ). We shall follow this approach in this proof, while in the meantime taking into account the intrinsic low-dimensional structure of the target distribution. In particular, we shall invoke the covering argument to treat the low-dimensional structure approximately as the union of linear subspaces. In particular, As it turns out, the evolution of the posterior mean function, i.e., $\mu_{t}(\cdot)$, can be well calibrated along the forward process with high probability under this framework, which plays a key role in enabling our non-asymptotic convergence analysis.
Our main proof is composed of three steps, with the first two steps being fairly standard and the third step capturing the main challenges. The analysis is supported by a sequence of auxiliary lemmas whose proofs are deferred to Appendices B and C. Throughout the proof, we adopt the following convenient notation:
@@ 5 Main analysis (Proof of Theorem 2)
In this step, we list several lemmas that help connect the KL error of interest with the mean squared difference between the true and estimated posterior means. These results, which have been used in past work like Chen et al., 2022b ; Chen et al., 2023a ; Benton et al., (2024); Azangulov et al., (2024); Wu et al., (2024); Potaptchik et al., (2024), are now standard for analyzing DDPM-type samplers.
@@
To cope with the term $T_{1}$, we find it helpful to first establish the following result on the coefficients $\{\eta_{t}\}$, whose proof is provided in Appendix B.1. The arguments of this step are also fairly standard, akin to the ones used in, e.g., Azangulov et al., (2024); Potaptchik et al., (2024).
@@
We now turn to the term $T_{2}$ in (21), which constitutes the main challenge of establishing linear-$k$ dependency. Here, $(Y_{t})_{t\in[0,T]}$ stands for the solution to the backward SDE (5) when initialized at $Y_{0}\sim X_{T}$, which satisfies $X_{t}\overset{\mathrm{d}}{=}Y_{T-t}$ (cf. (6)). Towards this end, we intend to first analyze each term in the integral. Define
@@
Towards this end, we proceed by characterizing the stochastic evolution of $\mu_{T-t}(Y_{t})$. We begin with the following lemma, which leverages Itô’s isometry to derive an alternative expression for $D_{s,t}$ that lends it well for subsequent analysis. The proof is provided in Appendix B.2.
@@
Remark 3.
Our approach shares similar spirits as Benton et al., (2024, Lemma 4). Similar characterizations were obtained in Section 4 of the concurrent work Potaptchik et al., (2024) by analyzing the martingale structure underlying the quantity of interest.
@@
We are now ready to present a key lemma that can assist in upper bounding $D_{s,t}$, which controls the size of the posterior covariance. The proof can be found in Appendix B.3.
@@
Remark 4.
Similar strategies were introduced in the previous work Benton et al., (2024, Appendix E), without accounting for the low-dimensional structure. It is worth noting that Lemma 9 in the concurrent work Potaptchik et al., (2024) shares similarity with the above result too. Note, however, that our lemma is established under much weaker assumptions on the low-dimensional structure as well as a much large range of the covariance of $X_{0}$, as discussed towards the end of Section 4.2.
@@ 6 Discussion
This paper has made progress towards demystifying the unreasonable effectiveness of then DDPM when the target distribution has a low-dimensional structure. We have sharpened the past result in Li and Yan, 2024a by proving that the iteration complexity scales linearly with the intrinsic dimension $k$ (note that linear dependency in $k$ was also established in the independent work Potaptchik et al., (2024), posted two weeks prior to our work). In some sense, our results reveal that the DDPM sampler is optimally adaptive to unknown low dimensionlity, provided that a suitable discretization schedule is adopted. Our theory leverages the intimate connection between the DDPM sampler and an adaptively discretized SDE, where the non-linear component of the drift time of this SDE is inherently low-dimensional.
We conclude this paper by pointing out several directions worthy of future investigation. To begin with, it remains unclear how the DDPM sampler adapts to unknown low dimensionality when other metrics (e.g., the total-variation distance or the Wasserstein distance) are used to measure distributional discrepancy. Another interesting direction is to examine the adaptivity of other mainstream samplers (e.g., the probability flow ODE (Song et al.,, 2021; Chen et al., 2023d, ; Li et al., 2024c, ; Gao and Zhu,, 2024; Huang et al.,, 2024)) in the face of unknown low-dimensional structure. Moreover, this paper treats the training procedure as a black box and directly assumes a mean squared score estimation error bound. For distributions with low intrinsic dimensions, a desirable score learning algorithm woulclud adapt to the low-dimensional structure too, which has been the focus of a recent line of work (Kadkhodaie et al.,, 2023; Tang and Yang,, 2024; Chen et al., 2024b, ; Azangulov et al.,, 2024; Chen et al., 2023b, ; Wang et al., 2024a, ; Wang et al., 2024b, ). It would thus be interesting to develop an end-to-end theory to accommodate both the pre-training phase and the sampling process. Finally, our analysis, which highlights the crucial role of proper parameterization and discretization, might shed light on how to design the coefficients for other algorithms (e.g., higher-order SDE and ODE solvers) in order to accommodate inherently structured data distributions.
@@ Acknowledgements
Y. Wei is supported in part by the NSF grant CCF-2418156 and the CAREER award DMS-2143215. Y. Chen is supported in part by the Alfred P. Sloan Research Fellowship, the AFOSR grant FA9550-22-1-0198, the ONR grant N00014-22-1-2354, and the NSF grant CCF-2221009. We would like to thank Jiadong Liang and Yuchen Wu for their inspiring discussions. We would also like to thank the authors of Potaptchik et al., (2024); Azangulov et al., (2024) for discussing the connections between our work and theirs, particularly for pointing out Azangulov et al., (2024, Proposition 28) that established the SDE equivalence of the DDPM.
Appendix A Proof of Theorem 1
Step 1: Solving SDE (11b).
Step 2: Establishing the equivalence between (36) and DDPM.
By defining
$\gamma_{n}=e^{-(T-t_{n})},$
we can explicitly calculate the integral as follows
Substitution into equation (37) then gives
Appendix B Proofs of the main lemmas
B.1 Proof of Lemma 6
@@ B.3 Proof of Lemma 8
where $C_{\mathsf{cover}}^{\prime}>0$ is a universal constant chosen as $C_{\mathsf{cover}}^{\prime}=C_{0}C_{\mathsf{cover}}$ (with $C_{0}$ specified in Assumption 1). Let $\{x_{i}^{*}\}_{1\leq i\leq s}$ be an $\varepsilon_{0}$-net of the data support $\mathcal{X}_{\mathsf{data}}$ (see the definition of an epsilon-net in Wainwright, (2019)), and let $\{\mathcal{B}_{i}\}_{1\leq i\leq s}$ be the corresponding $\varepsilon_{0}$-covering of $\mathcal{X}_{\mathsf{data}}$ obeying $x_{i}^{*}\in\mathcal{B}_{i}$ for all $1\leq i\leq s$. Without loss of generality, we choose the $\varepsilon_{0}$-covering in a way such that $\mathcal{B}_{i}$ and $\mathcal{B}_{j}$ are disjoint for all $1\leq i\neq j\leq s$, and $B(x_{i}^{*},\varepsilon_{0}/2)\subset\mathcal{B}_{i}$ for all $1\leq i\leq s$.
@@
$\mathcal{G}$$\coloneqq\big{\{}z\in\mathbb{R}^{d}\mid\|z\|_{2}\leq 2\sqrt{d}+\sqrt{C_{6}k\log k}\quad\text{and}$
@@
where $C_{6}>0$ is some sufficiently large constant. The index set $\mathcal{I}$ can be regarded as the collection of sets in $\varepsilon_{0}$-covering with at least reasonably large probability, and, as we will see momentarily, $\mathcal{G}$ represents a high-probability region of a standard Gaussian random vector. Further, define a high-probability region of the random vector $X_{t}$ as:
@@
whose high-probability property is confirmed by the following lemma. The proof of this lemma is provided in Appendix C.2.
@@
Next, to obtain tight control of $\mathbb{E}\left[\mathrm{Tr}\big{(}\Sigma_{t}(X_{t})\big{)}\right]$, we look at the events $\{X_{t}\in\mathcal{T}_{t}\}$ and $\{X_{t}\notin\mathcal{T}_{t}\}$ separately.
@@
For this case, the following lemma characterizes the typical performance of $X_{0}$, whose proof can be found in Appendix C.3.
@@
Consider any $x\in\mathcal{T}_{t}$, and suppose that $x=\sqrt{1-\sigma_{t}^{2}}x_{0}^{\prime}+\sigma_{t}z$ with $x_{0}^{\prime}\in\mathcal{B}_{i(x)}$ and $z\in\mathcal{G}$. Then
@@
holds for any $C\geq C_{7}$, where $C_{7}>0$ is some large enough constant.
Armed with the sub-Gaussian tail bound in Lemma 12, one can directly check that, for any $x\in\mathcal{T}_{t}$,
$\mathrm{Tr}\big{(}\Sigma_{t}(x)\big{)}\leq\mathbb{E}\big{[}\big{\|}X_{0}-x_{i(x)}^{*}\big{\|}_{2}^{2}\mid X_{t}=x\big{]}\lesssim\frac{\sigma_{t}^{2}}{1-\sigma_{t}^{2}}k\log k$ (47)
holds with $i(x)$ specified in Lemma 12.
@@
$\mathrm{Tr}\big{(}\Sigma_{t}(x)\big{)}\leq\mathbb{E}[\|X_{0}\|_{2}^{2}\mid X_{t}=x]\leq\sup_{x\in\mathcal{X}_{\mathsf{data}}}\|x\|_{2}^{2}\leq k^{2C_{R}}.$ (48)
@@
provided that $C_{6}$ is sufficiently large and $\sigma_{t}^{2}/(1-\sigma_{t}^{2})\geq\sigma_{\delta}^{2}\asymp\min\{\delta,1\}=\Omega(1/\mathrm{poly}(k))$. Inequality (49) taken together with inequality (44) concludes the proof.
Appendix C Proofs of the auxiliary lemmas
C.1 Proof of Lemma 1
@@
$N^{\mathsf{cover}}(\mathcal{M},\|\cdot\|_{2},\varepsilon_{0})\leq\frac{\mathsf{vol}(\mathcal{M})}{\omega_{k}}\left(\frac{\pi}{2}\right)^{k}\varepsilon_{0}^{-k},$
@@
$N^{\mathsf{cover}}(\mathcal{M}^{\varepsilon_{0}},\|\cdot\|_{2},\varepsilon_{0})\leq N^{\mathsf{cover}}(\mathcal{M},\|\cdot\|_{2},\varepsilon_{0}/2)\leq\frac{\mathsf{vol}(\mathcal{M})}{\omega_{k}}\pi^{k}\varepsilon_{0}^{-k}.$
@@
$\log N^{\mathsf{cover}}(\mathcal{M}^{\varepsilon_{0}},\|\cdot\|_{2},\varepsilon_{0})$ $\leq k\log\varepsilon_{0}^{-1}+\log\mathsf{vol}(\mathcal{M})-\log\omega_{k}+k\log\pi$
$\leq k\left(\log\varepsilon_{0}^{-1}+(\mathsf{vol}(\mathcal{M}))^{1/k}+C_{4}\right),$
where $C_{4}$ is some numerical constant. Therefore, if $\mathcal{M}$ satisfies $\tau_{\mathcal{M}}=\Omega(k^{-C_{1}})$ and $\mathsf{vol}(\mathcal{M})=O(k^{C_{2}})$ for some large enough constants $C_{1},C_{2}>0$, we arrive at
@@
C.2 Proof of Lemma 11
Since $\mathcal{T}_{t}$ is characterized through two different sets $\bigcup_{i\in\mathcal{I}}\mathcal{B}_{i}$ and $\mathcal{G}$, we analyze them separately.
Let us begin with the set $\bigcup_{i\in\mathcal{I}}\mathcal{B}_{i}$, which satisfies
$\mathbb{P}\bigg{(}X_{0}\notin\bigcup_{i\in\mathcal{I}}\mathcal{B}_{i}\bigg{)}$ $\overset{\mathrm{(a)}}{\leq}\sum_{i\in\mathcal{I}}\exp(-C_{6}k\log k)$